OpenStudy (anonymous):

I need help and I will give medals!!!!!!! find the surface area and the volume of three different three-dimensional figures. You will need to find three figures and measure the dimensions using a ruler or tape measure.

OpenStudy (perl):

how about a cube, a sphere, and a cylinder

OpenStudy (anonymous):

they will also work

OpenStudy (perl):

Cube : volume = s^3 , where s = side surface area = 6*s^2 , because there are 6 sides to a cube Sphere volume = 4/3*pi*r^3 surface area = 4*pi*r^2 cylinder Volume = pi*r^2*h surface area = 2pi*r*h + 2*pi*r^2

OpenStudy (anonymous):

can you solve the surface area and volume

OpenStudy (perl):

i wrote the formulas,